# Range Membership Inference Attacks

### v2

@@ Abstract
Machine learning models can leak private information about their training data, but the standard methods to measure this risk, based on membership inference attacks (MIAs), have a major limitation. They only check if a given data point exactly matches a training point, neglecting the potential of similar or partially overlapping data revealing the same private information. To address this issue, we introduce the class of range membership inference attacks (RaMIAs), testing if the model was trained on any data in a specified range (defined based on the semantics of privacy). We formulate the RaMIAs game and design a principled statistical test for its composite hypotheses. We show that RaMIAs can capture privacy loss more accurately and comprehensively than MIAs on various types of data, such as tabular, image, and language. RaMIA paves the way for a more comprehensive and meaningful privacy auditing of machine learning algorithms.
1 Introduction
@@
Machine learning models are prone to training data memorization [14, 15, 27, 42, 24]. It is also a known fact that the outstanding predictive performance of machine learning models on long-tailed data distributions often comes at the expense of blatant memorization of certain data points [15, 3, 30, 16]. In simple words, memorization is the phenomenon that models behave differently on training points, compared to other points. The memorization can lead to significant privacy risks as adversaries can infer private information about the training data from only having black box access to models.
@@
To quantify the privacy risk of machine learning models, a privacy notion needs to be agreed upon first. The reigning privacy notion is defined by membership information. Membership information of a data point is binary, but this single bit of information carries huge privacy implications. Being able to infer membership information opens up the possibility of conducting data reconstruction attack [37, 17, 4, 29], where the reconstruction attack inspects the membership of plausible data points to recover the training set. The de-facto way to audit the privacy risk according to this privacy notion is to conduct membership inference attacks (MIAs) [40], where an adversary aims to predict whether a given query data is part of the training set of the target model. The more powerful the membership inference attack is, the higher the privacy risk the target model bears.
Membership inference attacks provide a lower bound of the true privacy risk of a model, so improving the attack performance also means tightening the bound of privacy risk estimation. So far, the community has put much effort into improving the power of membership inference attacks by crafting better membership signals and constructing better statistical tests [39, 40, 36, 45, 5, 48]. While these have been useful for the betterment of privacy auditing, they have ignored the fundamental drawback of membership inference attacks as a practical privacy auditing tool, i.e., MIAs assume it is a privacy concern only if the adversary can identify the exact, full version of the training data. However, if the adversary can identify data points that are similar enough to the training points, it should also be treated as a significant privacy risk, because those points can contain similar levels of private information. For example, two Alice’s photos taken from slightly different angles, or with a different background would contain similar private information about Alice’s face or location. Similar to geometric transformations in images, small perturbations or rephrasing in textual data also affect little in the sensitivity of the information conveyed [13]. This oversight means private information leakage beyond exact matches of training data is currently ignored and the contemporary privacy auditing tools might produce overly optimistic results.
Besides, focusing on exact membership inference attacks renders them incapable of handling queries with missing values. This is another major limitation of MIAs, as data records with the same sensitive features and a few missing non-private features could carry a similar level of private information as the full data records. Imagine the case where an adversary can infer that an Asian person of age 25, identification number 123456 is in an HIV patients dataset, there is no need to identify or collect the exact values of the rest of the features, because the adversary is already able to pinpoint who has HIV from the given subset of key features. Even if the key identifiers are missing and unknown, it is still a grave privacy threat if the attacker can infer the rest of the features from the subset of known features that often contain quasi-identifiers, which has been extensively studied in the failure of k-anonymity as a privacy-preserving measure [38], where the attacker can reconstruct and identify training data with certain features masked. With existing inference attacks, there are two types of approaches to work with incomplete data. The first approach is to impute the missing features before performing membership inference attacks. However, if we simply make up the missing values or pass noisy data to the membership inference attack, the attack is expected to output “not a member,” since the chance of us coming up with the right values is extremely slim. If we use attribute inference [20] to help fill in the missing features, the imputed data may have inflated membership scores even if they are non-members, leading to a higher false positive rate. Therefore, existing MIA-based privacy auditing frameworks does not provide a natural solution to quantifying privacy leakage with missing features. The second approach that works with incomplete data is property inference [41, 1, 8, 9, 46], but it aims to extract statistical information about the features instead of testing membership information of data records. Hence, none of the existing inference attack frameworks can reliably quantify privacy leakage with incomplete data entries.
We argue that privacy quantification should not be point-based, because a small neighborhood around training points, defined based on a distance function out of the semantics of privacy, also contains similar private information. Hence, in this paper, we are proposing a new inference attack framework called range membership inference attacks (RaMIAs) to better capture the notion of privacy. Instead of using point queries and testing for exact matches, range membership inference attacks use range queries that cover a set of points. The goal of range membership inference attacks is to infer if the given range query contains any training point. In this paper, we denote the “neighborhood” as the “range”, and its corresponding distance function as the “range function”.
Range membership inference attacks extend the formulation of membership inference attacks. We adapt the original inference game formulation to reflect the change to range queries in RaMIAs. This extended formulation produces composite hypotheses in the likelihood ratio tests, which are the standard and best attack techniques in MIAs [39, 45, 5, 48]. Our method is based on standard statistical methods for composite hypothesis testing, namely generalized likelihood ratio tests (GLRTs) and Bayes factors. We show that RaMIAs can provide a more comprehensive notion of privacy by detecting private information leakage from the vicinity of training data when MIAs underestimate such privacy risk. Specifically, we observe a simple flipping can cause the membership score to decrease from very high to 0 (Figure 1), and the overall AUC can drop 20% if we test image classifiers with horizontally flipped images (Figure 2(c)). This is actually the expected behavior of MIAs since the transformed training data are non-members by definition. RaMIA, implemented with our simple attack algorithm (Sec 4), surpasses MIA by at least 5% on image datasets (Fig 3(b), 3(c)), providing better privacy auditing at the cost of as few as 15 samples, which is insignificant compared to the dimensionality of the data space.
In this paper, we emphasize the motivation and formulation of our newly proposed attack framework, RaMIA. As a proof-of-concept, we implement RaMIA with a simple attack strategy and experiment on tabular, image and text datasets, where RaMIA unanimously outperforms MIA. Additionally, our attack can also be potentially used in the pioneering membership inference and data extraction attacks on generative models [4, 44, 7], where the current evaluation requires finding the closest training image for all candidate data and computing their distances. By setting the distance function as the range function and conducting RaMIAs, we can evaluate the attacks more systematically.
@@
Figure 1: Examples of similar data with drastically different MIA scores. MIA would classify one of them as member, but the other as non-member. The dog image is from the CIFAR-10 dataset, while the text data is from AG News. The image classifer is not trained with horizontal flipping augmentation.
2 Preliminaries
2.1 Membership inference attacks
The membership inference attack (MIA) [40] is a type of inference attack against machine learning models to infer whether a given data sample is part of the model’s training set. Mathematically, given a model $f$ and a query point $\mathbf{x}$, the MIA aims to output 1 if $\mathbf{x}$ is a training point, and 0 otherwise. There are various methods to construct and conduct the attack, and it is still an active research direction that sees more powerful attacks being developed. Shokri et al. [40] use a shadow model-based approach where shadow models are trained on known training sets in similar ways to the target model. Confidence values of the training and test data on the shadow models are computed, which are then used as benchmarks in testing. However, the high cost and strong assumption of knowing the target model’s training details make the attack often infeasible. Yeom et al. [46] use model loss as a signal and threshold it, scraping the need for shadow models. Then MIA is formulated as an inference game (See Sec 3.1.1). Researchers turn to the principled approach to solve the game via likelihood ratio tests [39, 5, 45, 48]. Carlini et al. [5] and Ye et al. [45] propose reference-model based approaches, where target signals are compared to those obtained on reference models to obtain the likelihood ratio. To further boost the attack power, Zarifzadeh et al. [48] assumes the attacker has access to a pool of population data so that the likelihood ratio from reference-based attacks can be calibrated on ratios obtained on (non-member) population data.
@@ 2.1 Membership inference attacks
Recent attacks [5, 48] find extra attack performance on image data from augmenting the test queries with train-time augmentations, as these augmented training images could have been seen by the model during training. However, this assumes that the attacker knows the exact train-time augmentations in advance, and is able to sample from them. Augmenting training images with non train-time augmentations is not considered and allowed for valid reasons: those augmented images would be non-members in the current privacy notion based on membership inference, and testing them as members would violate the current privacy notion.
2.2 Range queries
If we make a connection to the field of databases, the existing membership inference attack operates on point queries or exact match queries. That is, each query to the membership inference attack only contains one data point and the attack only concerns if this very point is in the training set. On the other hand, range query, which is also a common querying operation in database systems, wants to retrieve all data points that fall into the ”range”. The most fundamental difference to point query is that the retrieved result often contains multiple data points instead of a single one. Our proposed attack, the range membership inference attack, operates with range queries.
3 From MIA to RaMIA
Membership inference attacks are often formulated as an inference game [46, 21, 45, 5, 48] between a challenger and an adversary. In this section, we will walk through how we come up with RaMIA from MIA.
3.1 Membership inference attacks
In membership inference attacks, the goal is to identify if a given point is part of the training set.
3.1.1 Membership inference game
@@
    The challenger samples a training dataset $D\overset{s_{D}}{\longleftarrow}\pi$, and trains a model $\theta\longleftarrow\mathcal{T}(D)$.
@@
    The challenger samples a data record $z_{0}\overset{s_{z_{0}}}{\longleftarrow}\pi$ from the data distribution, and a training data record $z_{1}\overset{s_{z_{1}}}{\longleftarrow}D$.
@@
3.1.2 Evaluation of MIA
Evaluation is done with a set of training and test points. True positive rate (TPR) and false positive rate (FPR) are computed by sweeping over all possible threshold values. By plotting the receiver operating characteristic curve (ROC), the power of an attack strategy can be represented by the area under the curve (AUC). A clueless adversary who can only randomly guess the membership labels is expected to get an AUC of 0.5. For stronger adversaries, they predict membership more accurately at each error level. Hence, they would achieve higher TPR at each FPR, and get a higher AUC.
3.1.3 Intrinsic limitation of MIA as a Privacy Auditing Framework
MIAs are intrinsically incapable of identifying points close to training points, regardless of how similar they are, because these points are, by definition, non-members in the scope of MIAs. Hence, there is a huge space of points that contain private information but are deemed non-members in the current privacy auditing framework. In this way, MIAs as privacy auditing tools become bad when the queries move away from the original data. Figure 2 shows the MIAs under-perform on non-original data. This inspires our formulation of RaMIA, where these points will be classified as ”members” for better and more comprehensive privacy auditing.
@@ 3.1.3 Intrinsic limitation of MIA as a Privacy Auditing Framework
Figure 2: MIA performance gets worse when the query points become further away from the training points. We define points different from training points but carry similar information as members. In 2(a), $m$ is the number of missing values. The query is constructed by filling in them with the most likely values. In 2(b), the point query changes to photos of the same identity who has at least one photo in the training set. In 2(c), the queries are horizontally flipped images. In 2(d), $d$ is the Hamming distance to original sentences.
3.2 Range membership inference attack
In range membership inference attacks, the goal is to identify if a given range contains any training point.
3.2.1 Range membership inference game
Here we define our range membership inference game, modified from the above formulation.
@@ 3.2.1 Range membership inference game
    The challenger samples a training dataset $D\overset{s_{D}}{\longleftarrow}\pi$, and trains a model $\theta\longleftarrow\mathcal{T}(D)$.
@@
    The challenger samples a data record $z_{0}\overset{s_{z_{0}}}{\longleftarrow}\pi$ from the data distribution, and a training data record $z_{1}\overset{s_{z_{1}}}{\longleftarrow}D$.
@@
    The challenger flips a fair coin to get the bit $b\in\{0,1\}$. If $b=1$, the challenger samples a range $\mathcal{R}_{1}$ containing at least one training point. Otherwise, challenger samples a range $\mathcal{R}_{0}$ containing no training points.
@@
The main difference between the two games is that the adversary now receives a range query (Step 4 in Def 2) instead of a point query (Step 3 in Def 1). We assume that the adversary is able to sample a set of points covered by any given range, which effectively means sampling within the range. This is a reasonable assumption because the adversary is usually assumed to have the ability to sample from the original data distribution $\pi$ [40, 45, 48] as specified in the membership inference game formulation. Given a range query and a sampler of $\pi$, it is not difficult to sample within the range.
@@
A range can be defined by a center, which is a point, a radius representing the size of the range, and a distance function which the radius is defined with. We refer to the center as the query center, the radius as the range size, and the distance function as the range function in this paper. One way to visualize a range is to imagine a unit $l_{2}$ ball around a point $x$, replacing the radius and $l_{2}$ distance with any arbitrary choice of range sizes and functions. Our framework can cater to any arbitrary range function that is potentially privacy-preserving. It can be spatially based (e.g. $l_{p}$ distances), transformation based (e.g. geometric transformations), and semantic based (e.g. owner/main features of the data). In the experiment section, we will present results with all of these types of range functions. Note that our attack reduces to user-level inference [31, 23, 27, 11, 10] when the range function is user-based, which means the privacy concern is whether any other data from the same owner are used in training. We will comment on the relations between RaMIA and other existing inference attacks later in the paper.
@@
In Step 3 of the range membership inference game, the details of how the challenger samples the ranges are intentionally omitted. This is because the ranges can be constructed around either in-distribution or out-of-distribution data points for both IN and OUT-ranges. The details of how we construct the iIN and OUT-ranges for our experiments are elaborated in Section 5.1.
3.3 Evaluation of RaMIA
Similar to MIAs, we evaluate RaMIA with AUCs. However, the notion of true positives and false positives are different from those defined in MIA, as both are defined on the range level. Specifically, if a range contains at least one training point, it is IN; otherwise, it is OUT. The TPR in RaMIA is the proportion of IN-ranges that are correctly identified by the attacker. On the other hand, the FPR in RaMIA is the proportion of OUT-ranges that yield positive predictions by the attacker. To avoid confusion, we call them (Range) TPR and (Range) FPR.
4 Range membership inference attacks
4.1 (Simple) Hypothesis testing

### v3

@@ Abstract
Machine learning models can leak private information about their training data. The standard methods to measure this privacy risk, based on membership inference attacks (MIAs), only check if a given data point exactly matches a training point, neglecting the potential of similar or partially overlapping memorized data revealing the same private information. To address this issue, we introduce the class of range membership inference attacks (RaMIAs), testing if the model was trained on any data in a specified range (defined based on the semantics of privacy). We formulate the RaMIAs game and design a principled statistical test for its composite hypotheses. We show that RaMIAs can capture privacy loss more accurately and comprehensively than MIAs on various types of data, such as tabular, image, and language. RaMIA paves the way for more comprehensive and meaningful privacy auditing of machine learning algorithms.
I Introduction
@@
Machine learning models are prone to training data memorization [1, 2, 3, 4, 5]. It is also a known fact that the outstanding predictive performance of machine learning models on long-tailed data distributions often comes at the expense of blatant memorization of certain data points [2, 6, 7, 8]. Memorization refers to the phenomenon where models behave differently on data depending on whether it was included in the training set. Such behavior can lead to significant privacy risks because adversaries can infer sensitive information about training data even with only black-box access to the model.
@@
To quantify the privacy risk of machine learning models, it is crucial to define a precise privacy notion. The prevailing privacy notion is based on membership information, a binary indicator that carries substantial privacy implications. Accurate inference of membership status can enable data reconstruction attacks [9, 10, 11, 12], where the adversary probes the membership of plausible data points to recover the training set. The de facto way to audit the privacy risk according to this privacy notion is to conduct membership inference attacks (MIAs) [13], where an adversary aims to predict whether a given point belongs to the training set of the target model. The more powerful the membership inference attack is, the higher the privacy risk the target model bears.
Figure 1: Illustration of privacy notions for MIAs and RaMIAs. LEFT: Sample CIFAR-10 training and test data are depicted with one highlighted IN range and one OUT range for visualization. In MIAs, only exact training points are deemed private, while all other points in the data space are considered non-private. The privacy notion is hence akin to a Dirac delta function that assigns privacy only to exact training points and zero everywhere else. RIGHT: RaMIA expands this definition by treating points that are similar to training data as privacy leaking. This approach captures privacy risks not only at exact training points but also in their vicinity, offering a more comprehensive and realistic privacy assessment.
Membership inference attacks provide a lower bound of the model’s true privacy risk, so improving the attack performance also tightens the bound of privacy risk estimation. So far, to more accurately audit the privacy risk, the community has largely focused on enhancing MIAs by developing stronger membership signals and more sophisticated statistical tests [14, 13, 15, 16, 17, 18]. While these advances have improved privacy auditing, they all quantify privacy risks by testing memorization of the exact, full version of training points. However, information leakage is not an all or nothing phenomenon. In many cases, a model may not memorize a training record in its entirety but rather partial information in the form of key features or data patterns. These memorization behaviors, which are more realistic and prevalent, lead to a form of leakage of sensitive information that can be exhibited from similar points in the vicinity of training data. For example, a model trained on images including people’s faces may capture the distinctive facial attributes that uniquely identify an individual while ignoring extraneous details such as the background. This could be a concern, as an image of the same person taken from a different angle or with a different background has a significant overlap in private information. Membership inference attacks, designed to detect exact matches, are not equipped to quantify this notion of information leakage as any slight change to training points would turn them into non-members where any correct MIA is expected to produce negative outcomes. For example, when using MIA, a simple horizontal flip can reduce a membership score from a high value to zero (Figure 2), and overall AUC can drop by 20% when testing image classifiers with horizontally flipped images (Figure 3(c)). This is expected as the transformed images are, by definition of MIA, non-members.
To address this gap, we propose a new class of inference attack, Range Membership Inference Attack (RaMIA), which is specifically designed to audit this new type of information leakage (Fig 1). The goal of RaMIA is to determine whether any training data exists within a defined neighborhood around a candidate point. Instead of relying on point queries that seek an exact match, RaMIA uses range queries that are defined by a center point, a distance measure that captures the semantics of privacy, and a radius that describes the area of interest. In practice, an auditor can craft a range query centered on any data record, image, or text. By applying a distance function that preserves the sensitive features (e.g., using the $\ell_{2}$ distance on unimportant features or tokens), the auditor can test whether the model has memorized these sensitive features even when the data tested are not exact replicas of any training data. In this way, privacy auditors can more accurately and comprehensively assess the privacy risk associated with sensitive information by tailoring the center, range function, and radius to address their specific concerns.
RaMIA is a flexible framework that supports a wide range of distance measures, enabling it to quantify various privacy risks based on different notions of proximity. Moreover, RaMIA aligns more closely with our intuitive understanding of privacy risk: when assessing the privacy of a data point, we expect the test to go beyond detecting exact matches and instead capture if any sensitive information is being memorized. For instance, if we claim that a model does not leak private information about my photo, we expect that it does not reveal any private information about me instead of merely the exact test photo. The range MIA enables formulating this notion of privacy risk.
RaMIA is also naturally related to current evaluation protocols for data extraction attacks on generative models [11, 21, 22], where candidate data points that fall within a specific tolerance are treated as successful extractions, reflecting the fact that close-by points can expose private information. Moreover, RaMIA can bypass naive test-time privacy protection schemes in which inputs that are recognized as training data are filtered (to fail naive use of MIA). Since the model disrupts the correspondence between inputs and exact training points, MIA is not expected to work properly. On the other hand, because RaMIA is capable of detecting information leakage in points close to training data, it can easily bypass this simple defense, thereby offering a more robust way of auditing information leakage against data reconstruction attacks.
Range membership inference attacks extend the formulation of exact membership inference attacks by incorporating range queries that capture privacy leakage in the vicinity of training data. We modify the traditional membership inference game to accommodate range queries, leading to composite hypotheses in the likelihood ratio tests, which are the standard and most effective techniques in MIAs [14, 16, 17, 18]. Our method leverages robust statistical approaches for composite hypothesis testing, including generalized likelihood ratio tests (GLRTs) and Bayes factors. We show that RaMIA provides a more comprehensive notion of privacy by detecting leakage from nearby training data when the direct application of MIAs underestimates the true risk. For example, testing horizontally flipped images instead of the original training and test images leads to a 20% drop in AUC when audited by MIAs (see Fig 3(c)). As a proof-of-concept, we implement RaMIA with a straightforward attack strategy and conduct experiments on tabular, image, and text datasets, where RaMIA consistently outperforms traditional MIA. Notably, even in the most challenging scenario where each IN range contains only one training point positioned near the boundary, our simple RaMIA algorithm (Sec IV) achieves a 5% improvement on image datasets (see Figs 4(b) and 4(c)). This result is obtained using at most 15 samples, which represents a negligible cost given the high dimensionality of the data space. It is important to note that these gains are observed when comparing RaMIA with the case where only traditional MIA is available and the MIA score of the center point is used to audit privacy with RaMIA’s membership definition. RaMIA’s performance further improves in less restrictive settings where each IN range contains multiple training points, underscoring its robustness and practical impact for comprehensive privacy auditing.
@@
Figure 2: Examples of similar data with drastically different MIA scores and potentially opposite membership predictions. The dog image is from the CIFAR-10 dataset, while the text data is from AG News. The image classifier is not trained with horizontal flipping augmentation.
II Related Work
II-A Membership inference attacks
The membership inference attack (MIA) [13] is a class of inference attack against machine learning models to infer whether a given data sample is part of the model’s training set. Mathematically, given a model $f$ and a query point $\mathbf{x}$, the MIA aims to output 1 if $\mathbf{x}$ is a training point, and 0 otherwise. Over time, various methods have been developed to construct and execute MIAs, making it an active research area with continuously evolving techniques. Shokri et al. [13] introduced a shadow model-based approach. In this method, multiple shadow models are trained on datasets that mimic the target model’s training set, and the confidence values for both training and test data are computed to serve as benchmarks. However, the high cost and strong assumption of knowing the target model’s training details make the attack often infeasible. [23] use model loss as a signal and threshold it, scraping the need for shadow models. Then MIA is formulated as an inference game (See Sec III-A1). Subsequent research has adopted a more principled approach by solving this inference game using likelihood ratio tests [14, 17, 16, 18]. [17] and [16] propose reference model-based approaches, where target signals are compared to those obtained on reference models to obtain the likelihood ratio. To further boost the attack power, [18] assumes the attacker has access to a pool of population data so that the likelihood ratio from reference-based attacks can be further calibrated on population data.
@@ II-A Membership inference attacks
Recent attacks [17, 18] find extra attack performance on image data from augmenting the test queries with train-time augmentations, as these augmented training images could have been seen by the model during training. This approach, however, presumes that the attacker knows the exact train-time augmentations and can replicate them. Augmenting training images with non train-time augmentations is deliberately excluded, as such images would be classified as non-members under the current privacy notion, and testing them as members would contradict the established definition.
Perturbation-based membership inference attacks
Several approaches have explored utilizing neighboring points around the query to improve MIA performance. [24] hypothesize that membership scores should remain consistently high in the vicinity of training points, and they aggregate scores from neighboring points to assess the smoothness of the scoring function. In contrast, [25] argue that the loss curvature would be sharper around training points, causing MIA scores to decline rapidly as one moves away from the training data. Although these perturbation-based methods incorporate neighboring information, they ultimately classify only the exact training points as members. Any perturbed or neighboring point is treated as non-member.
II-B Range queries
Drawing a parallel with database systems, traditional MIAs operate on point queries or exact match queries, where a single data point is retrieved. In contrast, a range query is designed to retrieve all data points within a specified interval or “range.” The key difference is that a range query often returns multiple data points rather than a single exact match. Our proposed attack, the range membership inference attack (RaMIA), builds upon this concept by operating with range queries, thereby extending the conventional MIA framework to capture privacy leakage in a broader context.
III From MIA to RaMIA
Membership inference attacks (MIAs) have traditionally been formulated as an inference game [23, 26, 16, 17, 18] between a challenger and an adversary. In this section, we first review the standard MIA framework and its evaluation protocol, then discuss its intrinsic limitations as a privacy auditing tool. These limitations motivate our proposed extension, range membership inference attacks (RaMIAs), which broaden the notion of membership to include points that leak sensitive information while not exactly in the training set, .
III-A Membership inference attacks
In membership inference attacks, the goal of the attacker is to identify if a given point is part of the training set.
III-A1 Membership inference game
@@
    The challenger samples a training dataset $D\longleftarrow\pi$, and trains a model $\theta\longleftarrow\mathcal{T}(D)$.
@@
    The challenger samples a data record $z_{0}\longleftarrow\pi$ from the data distribution, and a training data record $z_{1}\longleftarrow D$.
@@
III-A2 Evaluation of MIA
Conventionally, the MIA algorithm outputs a continuous MIA score for each point query. The membership decision is obtained by thresholding the score. Evaluation of MIAs is done on a set of training and test points. True positive rate (TPR) and false positive rate (FPR) are computed by sweeping over all threshold values. By plotting the receiver operating characteristic curve (ROC), the power of an attack strategy can be represented by the area under the curve (AUC). A clueless adversary who can only randomly guess the membership labels is expected to get an AUC of $0.5$. Stronger adversaries predict membership more accurately at each error level. Hence, they would achieve higher TPR at each FPR, and get a higher AUC.
III-A3 Intrinsic limitation of MIA as a Privacy Auditing Framework
MIAs are designed to be incapable of identifying points close to training points, regardless of how similar they are, due to the strict definition of members. Hence, there is a vast data space of points that contain private information but are deemed to be non-members in the current privacy auditing framework. In this way, MIAs as privacy auditing tools become out-of-scope when the queries move away from the original data, resulting in unpredictable and unreliable auditing results in these scenarios. Figure 3 illustrates how MIA performance deteriorates as the query points deviate from the original training data. This observation motivates our formulation of RaMIA, where we broaden the definition of membership to capture privacy leakage from similar data points.
@@ III-A3 Intrinsic limitation of MIA as a Privacy Auditing Framework
Figure 3: MIA performance gets worse when the query points become further away from the training points. We define points different from training points but carry similar information as members. In 3(a), $m$ is the number of missing values. The query is constructed by filling in them with the most likely values. In 3(b), the point query changes to photos of the same identity who has at least one photo in the training set. In 3(c), the queries are horizontally flipped images. In 3(d), $d$ is the Hamming distance to original sentences.
III-B Range membership inference attack
In range membership inference attacks, the goal is to identify if a given range contains any training point. We define our range membership inference game, modified from the MI game.
@@ III-B Range membership inference attack
    The challenger samples a training dataset $D\longleftarrow\pi$, and trains a model $\theta\longleftarrow\mathcal{T}(D)$.
@@
    The challenger samples a data record $z_{0}\longleftarrow\pi$ from the data distribution, and a training data record $z_{1}\longleftarrow D$.
@@
    The challenger flips a fair coin to get the bit $b\in\{0,1\}$. If $b=1$, the challenger samples a range $\mathcal{R}_{1}$ containing $z_{1}$. Otherwise, challenger samples a range $\mathcal{R}_{0}$ containing $z_{0}$ and no training points.
@@
The key difference is that the adversary now receives a range query (Step 4) rather than a single data point. We assume that the adversary can sample a set of points from any given range—a reasonable assumption given their ability to sample from the data distribution $\pi$ as in traditional MIAs [13, 16, 18].
@@
A range can be defined by a center, which is a point, a radius representing the size of the range, and a distance function which the radius is defined with. We refer to the center as the query center, the radius as the range size, and the distance function as the range function in this paper. Formally, we can define a range by =$\mathcal{R}=\{x^{\prime}:d(x^{\prime},x)\leq\epsilon\}$, where $x$ is the range center, $d$ is the range function and $\epsilon$ being the range size. One way to visualize a range is to imagine an $l_{2}$ ball around a point $x$, replacing the radius and $l_{2}$ distance with any arbitrary choice of range sizes and functions. Our framework is flexible to accommodate any distance function that preserves a significant amount of the sensitive information. The range function can be spatial (e.g. $l_{p}$ norms), transformation-based (e.g. geometric transformations), or semantic (e.g., based on user identity). In the experiment section, we will present results with all of these types of range functions. Notably, RaMIA reduces to user-level inference [27, 28, 3, 29, 30] when the range function is defined on a per-user basis.
@@
In Step 3 of the range membership inference game, the specific procedure for constructing ranges is intentionally left unspecified. This is for flexibility: ranges can be constructed around either in-distribution (ID) or out-of-distribution (OOD) data points for both IN and OUT cases. The details of our range construction methods for experiments are provided in Section V-A.
III-C Evaluation of RaMIA
RaMIA is evaluated similarly to MIAs using AUC metrics, but with definitions adapted to the range setting. A range is considered IN if it contains at least one training point and OUT otherwise. Thus, the TPR is defined as the proportion of IN ranges correctly identified by the adversary, and the FPR is the proportion of OUT ranges incorrectly classified as IN. To avoid confusion, we call them (Range) TPR/FPR.
IV Range membership inference attacks
IV-A (Simple) Hypothesis testing